Suppose that the following information describes the economy of Wonderland: C = 300 + 0.8Yd I = 300 G = 250 X = 300 M = 150 + 0.6Yd T= 0.25Yd Wonderland’s multiplier is calculated as:


1
Expert's answer
2021-11-14T17:36:47-0500

In an open economy, when talking about multiplier, one need to know that:

mps+mpc+mpm+mpt=1

Where,

  • mps is marginal propensity to save
  • mpc is marginal propensity to consume
  • mpt is marginal propensity to tax
  • mpm is marginal propensity to imports

According to the given information, Wonderland’s value of multiplier would be:

"\frac{1}{MPC}=\frac{1}{1-0.8}=5\\or\\\frac{1}{MPs+mom+mpt}=\frac{1}{-0.65+0.6+0.25}\\\frac{1}{0.2}=5"